write the equation of the line perpendicular to 2x-3y=7 through the point (-6,5)
Answer by Greenfinch(383)

3y = 2x -7 or y = (2/3)x - 7/3
Gradient is 2/3 (M m = -1)
Line perpendicular to this is thus -3/2
So is 2y + 3x = c
Going through (-6,5) gives -12 + 15 = c
So c = 3
Equation is therefore
2y + 3x = 3